A debt of x, with interest rate of 7% compounded annually will be retired at the end of 10 years through the accumulation of deposit in the sinking fund invested at 6% compounded semi-annually. The deposit in the sinking fund every end of six months is 21,962.68. What is the value of x?
Answer by Theo(13342)

the future value of your investment of 21,962.68 at the end of every 6 month period for 10 years is equal to $590,145.44, according to the results of the financial calculator i used at https://arachnoid.com/finance/

that is the just enough to pay off your debt at that time.

the future value of your debt at 7% per year compounded anually is equal to x * 1.07^10 = 1.967151357 * x.

therefore 1.967151357 * x = 590,145.44.

solve for x to get x = 590,145.44 / 1.967151357 = 300,000.0167.

this can be rounded to 300,000 dollars, which should be your answer.

the future value of your debt of 300,000.0167 today is equal to 300,000.0167 * 1.07 ^ 10 = 590145.44.

the future value of your semi-annual payments into the sinking fund is exactly equal to that as shown in the following printout.

$$$

the inputs to that calculator were:

present value = 0
number of time periods = 10 years * 2 payments per year = 20 semi-annual time periods.
the payment made at the end of each semi-annual period is 21962.68.
the interest rate per time period is 6% per year / 2 semi-annual periods per year = 3% per semi-annual time period.
the payments are made at the end of each semi-annual time period.

the output of that calculator is:

future value = 590,145.44.